Project

General

Profile

NDOS Commissioning Meeting 2011/05/19

Attendance: Jon, Gary, Susan, Zeb, Jan, Leon, Minerba, Mark Messier, Mark Mathis, Peter, Gavin, Denis, Mat, Ken

Agenda:

  1. NDOS Status and Channel Recovery Plans
  2. DAQ/Electronics Update & Plans
  3. DCS Update & Plans
  4. DataCheck Update
  5. AOB (All)

NDOS Status

  • Jon:
    • uptime (run>30 minutes long and all 9 dcms in the readout) during the last week was ~70%
    • active channel fraction (whole detector) dropped from ~78% to 65%. The following recent changes were observed in each DCM:
      • dcm-3-04-02: ~90% --> ~56%
      • dcm-3-04-01: ~97% --> ~80%
      • dcm-3-03-01: ~97% --> ~85%
      • dcm-3-02-03: ~72% --> ~42%
      • dcm-3-02-02: ~70% --> ~42%
      • dcm-3-02-01: ~85% --> ~73%
      • dcm-3-01-03: ~90% --> ~85%
      • dcm-3-01-02: ~95% --> ~91%
      • dcm-3-01-01: ~87% --> ~80%
    • Denis: some of the losses are due to cooling, where APDs are lost permanently after cooling
    • Mat: humidity in the hall is high again (50%). Jon will contact Bill Moorehouse about this. Discussion about why electronics are sensitive to humidity. Leon would like to try to run a FEB in an environmental chamber, but it will take some time to develop the test.
  • Mat: installed 15 cable-ectomized cables on side of di-block 2
    • details are in doc-db 5980
    • only 2/15 show an improvement; further study/thought is needed
  • Denis: cooled DCM 4-2. Found a few FEBs had water lines swapped. Found a water leak (leaking on connection to heatsink), entire FEB was wet, so it was turned off.

DAQ/Electronics Update & Plans

  • Peter:
    • lots of stress testing (config. parameters set to value well beyond the nominal settings). Andrew N. has been leading the charge.
    • corrupted data is now being seen, but it's not clear if it's due to the stress-testing. Bug was found and fixed in BNEVB; data corruption rate has decreased but not entirely.
  • Mark suggests making raw data unpacking code more robust.
  • Peter: could be version skew in Dispatcher.
  • Peter:
    • BNEVB was found to not to be able to handle high trigger rates, but this was expected and it is understood how to fix.
    • dcm "names" have been changed to their logical names, not hardware names.
    • Gennadiy made change to DDS that has improved stability of the messaging system.

DCS Update & Plans

  • Leon: bug with DCM firmware that was causing it to lock when external<->internal command switch is made is being worked on.

DataCheck Update

AOB (All)